Let individual i have the utility function

Ui(x,y) = x^1/2 * y

Suppose the initial endowment is x=16 and y =2.

A) Determine 4 combinations (x,y) for which Ui=8, and for Ui-16

B) What is the minimum amount of x that i would have to be compensated to induce him to give up 1 unit of y?
Answer #1

Ui = x1/2y

A) Ui = x1/2y
So, x1/2 = Ui/y
So, x = (Ui/y)2

Ui = 8
When y = 1; x = (8/1)2 = 64; (x,y) = (64, 1)
When y = 2; x = (8/2)2 = 16; (x,y) = (16, 2)
When y = 4; x = (8/4)2 = 4; (x,y) = (4, 4)
When y = 8; x = (8/8)2 = 1; (x,y) = (1, 8)

Ui = 16
When y = 1; x = (16/1)2 = 256; (x,y) = (256, 1)
When y = 2; x = (16/2)2 = 64; (x,y) = (64, 2)
When y = 4; x = (16/4)2 = 16; (x,y) = (16, 14)
When y = 8; x = (16/8)2 = 4; (x,y) = (4, 8)

B) MRSyx = MUy/MU
MUx =

MUy =

So, MRSyx = x1/2/(1/2)x-1/2y = 2x(1/2)+(1/2)/y = 2x/y
At x = 16, and y = 2
MRSyx = 2*16/2 = 16
So, the minimum amount of x that i would have to be compensated to induce him to give up 1 unit of y is 16 units.